OpenStudy (anonymous):

cot(108^0) A) 2.47 B) -0.32 C) 0.41 D) -3.08 i need to learn how to do the cot

OpenStudy (anonymous):

cotx=1/tan x

OpenStudy (campbell_st):

find tan(108) then take the reciprocal

OpenStudy (zzr0ck3r):

= cos(108)/sin(108)

OpenStudy (campbell_st):

its 2nd quadrant so will be negative

OpenStudy (anonymous):

i thought 1/tan = cos/sin @Krishnadas

OpenStudy (campbell_st):

tan = sin/cos so the reciprocals are are you have written

OpenStudy (anonymous):

@arria both are same

OpenStudy (anonymous):

tanx=sinx/cosx ---->cotx=1/tanx=cosx/sinx